OpenStudy (anonymous):

if 50ml each of nitric oxide and oxygen is mixed , the volume of nitrogen dioxide produced is ?

OpenStudy (anonymous):

first you need to write chemical reaction equation, then calculate moles of reactants and from ratio of product and one of the reactants and amount of one of the reactants calculate moles of reactant that are created and from amount calculate volume n=V/Vm

OpenStudy (anonymous):

can u explain n=V/Vm

OpenStudy (anonymous):

n - amount of reactant/product in moles V - volume Vm - molar volume - constant - 22,4 dm3mol-1

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

no problem, if you need further explanation just say what is unclear...

OpenStudy (anonymous):

hmmm.. do u mind solving the problem and showing the steps..?so that i can understand better

OpenStudy (anonymous):

NO + 1/2 O2 -> NO2 n(NO)= 50 * 10^-3 dm3 / 22,4 dm3mol-1 = 2,23mol n(O2) = 2,23 mol n(NO2)= n(NO)=2,23 V(NO2)=n(NO2)*22,4 dm3mol-1= 50 ml

OpenStudy (anonymous):

if you have more questions ask them but i will answer later cause i have to go to lab now my students wont wait :)

OpenStudy (anonymous):

k ..thank you .. :)